Calculate the critical angle for crown glass surrounded by water. The index of refraction of crown glass is 1.52 and for water, it's 1.333.
Answer #1

Using snell's law:

n1*sin A1 = n2*sin A2

n1 = refractive index of water = 1.333

when A1 = 90 deg, Critical angle will be

n2*sin Ac = n1*sin 90 deg

Ac = arcsin (n1/n2)

refractive index of crown glass = 1.52

Ac = arcsin (1.333/1.52) = 61.28 deg

Please Upvote.